$5400 is invested, part of it at 10% and part of it at 7%. For a certain year the total yield is $528. How much was invested at each rate?
Answer by sachi(548)

let $x is invested at 10% and $(5400-x)is it at 7%
so the total yield is
10x/100+7(5400-x)/100=528
or 10x+37800-7x=52800 multiplyig 100 on both sides
or 3x=52800-37800=15000
or x=5000
so $5000 is invested at 10% and $400 is it at 7%
ans